OpenStudy (anonymous):

Super stuck on this.. any help would be great! Thanks Determine the solution of the system. y = x + 6 y = 2x

jimthompson5910 (jim_thompson5910):

The idea here is to plug in y = 2x into the first equation, solve for x, then use it to find y. y = x + 6 2x = x + 6 2x-x = 6 x = 6 ----------------- y = x+6 y = 6+6 y = 12 So the solutions are x = 6 and y = 12 giving us the ordered pair (6,12)

OpenStudy (anonymous):

Wow that is a lot easier than I thought. I guess I was just over thinking it. Thank you again! You have been very helpful!

jimthompson5910 (jim_thompson5910):

you're welcome :)